THE EFFECTS OF TRIPLET PERTURBERS ON PHOTOPHYSICAL PROCESSES IN ACETYLENE

\underline{Stephen Drucker}, Jonathan P. O’Brien, Paresh Patel, and Robert W. Field, MIT, 77 Massachusetts Avenue, Cambridge, MA 02139.

We have recorded excitation spectra of the acetylene Ã^{1}A_{u} ← X̃^{1}Σ_{g}^{+} transition by measuring simultaneously the total emission in two spectral regions: UV (200-400 nm) and NIR (0.9-1.5 μm). UV emission is due to well-characterized^{1} Ã → X̃ transitions. The source of the NIR signal is suggested by the emission patterns within the subbands studied, V_{0}^{2}K_{0}^{1} and V_{0}^{3}K_{0}^{1}. (V denotes the trans-bending mode.) The NIR:UV intensity ratio is approximately constant throughout the V_{0}^{2}K_{0}^{1} subband. This is consistent with excited C_{2}H or C_{2} as the carrier of the NIR signal, because these species, capable of emitting in the NIR^{2,3}, could be produced via multiphoton processes^{2} resonantly enhanced by the Ã state as intermediate. The V_{0}^{3}K_{0}^{1} subband shows significantly more fluctuation in the NIR:UV intensity ratio than does the V_{0}^{2}K_{0}^{1} subband. The one-quantum increase in trans-bending excitation enables accidental resonances with background states to modulate the NIR emission pattern. Triplet perturbers could participate in this process by modifying the resonance enhancement of C_{2}H or C_{2} production. Mixed S_{1}–triplet states could additionally contribute to the NIR signal through T_{n} → T_{1} (n > 1) transitions.

1K. Yamanouchi, N. Ikeda, S. Tsuchiya, D. M. Jonas, J. K. Lundberg, G. W. Adamson, and R. W. Field, J. Chem. Phys. 95, 6330 (1991). 2Y.-C. Hsu, M.-S. Lin, and C.-P. Hsu, J. Chem. Phys. 94, 7832 (1991). 3F. Shokoohi, T. A. Watson, H. Reisler, F. Kong, A. M. Renlund, and C. Wittig, J. Phys. Chem. 90, 5695